Markthegoat1117:

Solve: 3 + 2 • (8 – 3)

fendi:

13

Kagura:

brand new me

Kagura:

Ans. 3 + 2 • (8 – 3) = 3 + 2 (5) = 3 + 10 = 13

Kagura:

GO CRAZYYYY

fendi:

ahhh

Kagura:

i love that song

fendi:

smd

Kagura:

no u

fendi:

i got scared at first

Kagura:

LMAO why

fendi:

my screen just turned white

JayLover:

13